1. Suppose that the economy of Loserville has the following distribution of income:

Quintile                                   share of income   

            first quintile     (20%)              10% of aggregate income                  

            second quintile (40%)             15% of aggregate income                  

            third quintile    (60%)             20% of aggregate income                  

            fourth quintile (80%)             25% of aggregate income                  

            fifth quintile    (100%)           30% of aggregate income                

What is the Gini Cofficient (or Index) for Loserville?

a.

.15 (or 15)

b.

.40 (0r 40)

c.

.57 (or 57)

d.

.20 (or 20)

e.

.30 (or 30)

2.

If a country has a surplus in its current account

a.

then it must have a deficit in its capital and financial accounts.

b.

then its net exports are positive.

c.

then it is selling more to the rest of the world than it is buying from the rest of the world.

d.

then the country has a trade surplus.

e.

all of the above.

3.

Suppose that while sitting around bored in your dorm room one night, you and your friends decide to construct an Easter Basket Price Index (EBPI). After careful research, you come up with the following data:

           

item

quantity

price(2019)

price(2020)

basket

1

$4.00

$5.00

grass

2

$1.00

$2.00

cream eggs

2

$2.00

$3.00

jelly beans

4

$0.50

$1.00

bunnies

2

$4.00

$6.00

chicks

4

$2.00

$3.00

eggs

6

$1.00

$2.00

stuffed bunny

1

$10.00

$15.00

Hershey's kisses

10

$0.60

$1.00

Assume that 2019 is the base year.

What is the Easter Basket Price Index for 2020?

a.

100

b.

50

c.

80

d.

160

e.

60%
Answer #1

1.

Area of 1 = 0.5 * base * height = 0.5 * 20 * 10 = 100

Area of 2 = 0.5 * base * height = 0.5 * (40-20) * (25 - 10) = 0.5 * 20 * 15 = 150

Area of 3 = 0.5 * base * height = 0.5 * (60-40) * (45 - 25) = 0.5 * 20 * 20 = 200

Area of 4 = 0.5 * base * height = 0.5 * (80-60) * (70-45) = 0.5 * 20 * 25 = 250

Area of 5 = 0.5 * base * height = 0.5 * (100-80) * (100-70) = 0.5 * 20 * 30 = 300

Area of 6 = Length * Width = (100-80) * (70-45) = 20 * 25 = 500

Area of 7 = Length * Width = (100-60) * (45-25) = 40 * 20 = 800

Area of 8 = Length * Width = (100-40) * (25-10) = 60 * 15 = 900

Area of 9 = Length * Width = (100-20) * (10-0) = 80 * 10 = 800

Total Area under the Lorenz curve = Areas of (1+2+3+4+5+6+7+8+9) = 4000

Total Area under line of total equality (45o line) = 0.5 * 100 * 100 = 5000

Area between the line of total equality and the Lorenz curve = 5000 - 4000 = 1000

GINI Coefficient = Area between the line of total equality and the Lorenz curve/Total Area under line of total equality = 1000/5000 = 0.20

Ans: d. 0.20 (or 20)